# Visual Planning: Let’s Think Only with Images

### v2

@@ C.2.1 Examples of trained Textual variants
C.3 Performance with Scaling Difficulties
@@ C.4 Out-of-Distribution Performance
C.5 Ablation: The Role of Stage 1
@@
C.6 Visual Planning Results
VPRL Stage 1 and Stage 2
Table 10 presents results for each stage of VPRL. After Stage 1, the model learns to generate plausible images but lacks goal-directed behavior, resulting in near-random performance across tasks. In Stage 2, reinforcement learning instills purposeful planning, enabling the model to align generations with the goal and outperform VPFT across all benchmarks.
@@ C.6 Visual Planning Results
Generated Visual Planning Traces for Illustration
Figure 13 shows the generated visual planning traces for FrozenLake, with Figure 14 for Maze and Figure 15 for MiniBehavior. Each visual trajectory begins with the initial state as the input (the first frame), followed by a sequence of intermediate states generated by VPRL that form the predicted visual plan.
@@
Appendix D Prompting Templates

### v3

@@ F.2.1 Examples of trained Textual variants
F.3 Performance with Scaling Difficulties
@@ F.4 Out-of-Distribution Performance
F.5 Ablation: The Role of Stage 1
@@
F.6 Visual Planning Results
VPRL Stage 1 and Stage 2. Table 10 presents results for each stage of VPRL. After Stage 1, the model learns to generate plausible images but lacks goal-directed behavior, resulting in near-random performance across tasks. In Stage 2, reinforcement learning instills purposeful planning, enabling the model to align generations with the goal and outperform VPFT across all benchmarks.
@@ F.6 Visual Planning Results
Generated Visual Planning Traces for Illustration. Figure 13 shows the generated visual planning traces for FrozenLake, with Figure 14 for Maze and Figure 15 for MiniBehavior. Each visual trajectory begins with the initial state as the input (the first frame), followed by a sequence of intermediate states generated by VPRL that form the predicted visual plan.
@@
F.7 Image Quality Analysis
Figure 16: Qualitative comparison between original images (top), predicted images by the model (middle), and reconstructed images obtained by encoding and decoding the original inputs (bottom).
Table 11: Exact Match (EM) and Progress Rate (PR) on FrozenLake under VPRL when using ground-truth images versus self-generated images as inputs during inference.
Model EM (%) PR (%) L3 L4 L5 L6 Avg. L3 L4 L5 L6 Avg. VPRL   xxx- (w/ self-generated images) 97.6 95.6 90.8 82.4 91.6 98.4 96.0 93.0 85.6 93.2   xxx- (w/ ground-truth images) 98.4 95.2 93.2 81.6 92.1 98.5 95.8 94.1 85.3 93.4
It can be observed that the intermediate images on FrozenLake generated by the visual planner in Figure 4 contain noticeable artifacts, and we suspect that this noise arises from the limitation of the image tokenizer rather than from the model’s image generation ability. To verify this, we include an additional analysis on FrozenLake that illustrates how the tokenizer reconstructs images in our framework.
Limitations of the Image Tokenizer. Figure 16 confirms that the artifacts observed in our predicted images originate from the tokenizer rather than from the prediction process itself. When encoding a ground-truth image into visual tokens and decoding it back, the reconstructed output shows similar artifacts inevitably introduced by the tokenizer to those in the model’s predictions, which makes the reconstruction not identical to the original image. At the same time, we observe that the intermediate images produced by the model are already comparable in quality to the reconstructed images. While our work focuses on planning rather than image generation quality, this observation indicates that the visual planner generates images that are sufficient for effective planning.
We consider this behavior to be encouraged by the dynamics interpreter. During the training, the dynamics interpreter serves as an implicit format constraint. Any generated image that it cannot parse is treated as an invalid transition and receives a penalty, enforcing the model to maintain the semantic structure of the environment in its generated images.
Robustness to Intermediate Image. We subsequently conduct a quantitative study to evaluate whether providing high-quality intermediate images at inference improves performance. Instead of feeding back the model’s self-generated image at each step, we replace it with the ground-truth image rendered by the environment, which serves as a high-quality version.
Table 11 shows that the performance with and without high-quality images remains similar across all grid sizes. This shows that our visual planner is robust to visual noise and does not depend on perfectly rendered images to plan effectively.
F.8 Computational Cost Analysis
Table 12: Average inference token cost across FrozenLake, Maze, and MiniBehavior. We also report the average of the task-level average costs. Higher values indicate higher computational cost.
Model FrozenLake Maze MiniBehavior Avg.
Closed-Source Models
Gemini 2.0 Flash
  xxx- Direct 10.8 12.5 14.8 12.7
  xxx- CoT 150.5 166.5 196.5 171.2
Gemini 2.5 Pro (think) 885.6 1030.2 1619.9 1178.6
Open-Source Models
Qwen 2.5-VL-Instruct-7B
  xxx- Direct 13.4 95.9 13.9 41.1
  xxx- CoT 306.2 316.4 272.3 298.3
  xxx- SFT 10.7 11.4 13.2 11.8
LVM-7B
  xxx- VPFT (ours) 819.2 957.2 1471.2 1082.5
  xxx- VPRL (ours) 819.2 957.2 1471.2 1082.5
Table 13: Average inference token cost of trained textual planner variants on FrozenLake.
Model Token Cost
Qwen 2.5-VL-Instruct-7B
- SFT
  xxx- Direct 10.7
  xxx- w/ Coordinates 179.0
  xxx- w/ ASCII 84.3
- GRPO
  xxx- w/ VPRL progress reward 129.8
  xxx- w/ PR metric reward 74.9
To provide a quantitative comparison of the computational cost between visual planning and traditional textual reasoning, we further analyse the token usage of both the visual planner and the textual baselines during inference. We compute the average number of generated tokens for all models reported in Table 1 across all tasks. In addition, we include a more detailed breakdown of the token cost for the trained textual planner variants listed in Table 2, evaluated on FrozenLake.
Table 12 and Table 13 summarise the resulting inference token cost. As expected, visual planning introduces a noticeable computational overhead due to repeated image generation. However, this additional cost remains affordable in practice when compared with textual CoT. On average across the three tasks, the token cost of our visual planner is roughly 3 times that of Qwen 2.5-VL-Instruct-7B with CoT and around 6 times that of Gemini 2.0 Flash with CoT, suggesting that our method is still computationally feasible. We also observe that thinking models, such as Gemini 2.5 Pro, produce the largest number of tokens among all tasks, indicating that visual planning is not always the most expensive option.
Appendix G Prompting Templates